# Towards the understanding of heavy quarks hadronization: from leptonic to heavy-ion collisions

### v2

@@ III.4 Measurements as a function of multiplicity
The LHCb Collaboration recently released the measurements of the $\sigma_{\Lambda_{\mathrm{b}}^{0}}/\sigma_{\mathrm{B}^{0}}$ as a function of $p_{\mathrm{T}}$ and multiplicity [194]. The $p_{\mathrm{T}}$-integrated $\sigma_{\Lambda_{\mathrm{b}}^{0}}/\sigma_{\mathrm{B}^{0}}$ ratio is shown as a function of relative multiplicity in Fig. 20. In the lowest multiplicity interval, the LHCb data approaches the baryon fraction measured in $\mathrm{e^{+}e^{-}}$ collisions at the LEP. The distinct rise of the baryon fraction with multiplicity reaches a plateau at $\sim$ 0.5 for collisions that produce more than twice the average number of VELO tracks.
These last two measurements show that b-quark hadronization is different in $\mathrm{e^{+}e^{-}}$ and pp collisions, thus it is not an universal process. This may provide indication that quark coalescence plays an important role in forming beauty hadrons at the LHC, or that one should more carefully examine string models with multiplicity dependencies. The trend observed for the $p_{\mathrm{T}}$-integrated $\sigma_{\Lambda_{\mathrm{b}}^{0}}/\sigma_{\mathrm{B}^{0}}$ ratio resembles a “turn-on” curve which may mark the transition from two different regimes, with the onset of a different hadronization mechanism in pp collisions becoming very quickly effective as soon as a minimal hadronic activity is present. It would be interesting to try to correlate the observed trend with the number of MPI. It must be noted that the evolution observed for the $p_{\mathrm{T}}$-integrated $\sigma_{\Lambda_{\mathrm{b}}^{0}}/\sigma_{\mathrm{B}^{0}}$ ratio does not contrast with what previously discussed about the evolution with multiplicity of the $p_{\mathrm{T}}$-differential $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio measured by ALICE at midrapidity. This will be further discussed in Sect. III.6.
@@ III.5 Proton–nucleus collisions
The LHCb collaboration measured the $\Lambda_{\mathrm{c}}^{+}$ and $\Xi_{\mathrm{c}}^{+}$ production and their ratios to $\mathrm{D^{0}}$ at forward and backward rapidity in p–Pb collisions [218, 219]. When compared with the measurements at midrapidity [220, 173], the baryon-to-meson ratios are lower for $p_{\mathrm{T}}$ $<$ 8 GeV/$c$, while at higher $p_{\mathrm{T}}$ the measurements are consistent within uncertainties. The origin of this tension is not yet fully understood, for a more detailed discussion on this open point see Sect. IV.
@@ III.6 Nucleus–nucleus collisions
Several measurements in Au–Au collisions at RHIC and Pb–Pb collisions at the LHC demonstrate that charm and beauty quarks interact strongly with the QGP constituents [224, 225, 226, 227, 228, 161, 229, 230, 231, 232, 233, 234, 235, 236, 237, 238, 239, 240, 241, 242, 243, 244, 245, 246, 247, 159, 248, 249, 250, 251, 252]. The $R_{\mathrm{AA}}$ values observed for both charm and beauty hadrons are significantly smaller than unity at intermediate $p_{\mathrm{T}}$ and larger for beauty than charm, indicating a significant energy loss of both quarks in the medium and evidencing the expected dependence of energy loss on the quark mass. At moderate/low $p_{\mathrm{T}}$ the $v_{2}$ of prompt D mesons is large, close to that observed for charged pions [237]. This can indicate that charm quarks are strongly coupled to the system and are pushed towards kinetic equilibrium via rescattering with the medium constituents. It can also signal coalescence as a hadronization process relevant to charm quarks through which charm hadrons partly inherit the flow of light quarks, a relevant aspect that could be more deeply investigated through heavy-light correlation of the anisotropic flow $v_{n}$ [253, 84]. At low $p_{\mathrm{T}}$, where elastic processes are the most important ones, heavy quark transport in the medium can be described as a diffusion process [254, 255, 55, 256]. The main transport parameter is the spatial diffusion coefficient $D_{s}$, which is related to the kinetic equilibration time of a heavy-quark with mass $m_{\mathrm{Q}}$, $\tau_{\mathrm{Q}}=D_{s}m_{\mathrm{Q}}/T$ [257, 258]. The comparison of $R_{\mathrm{AA}}$ and $v_{\mathrm{2}}$ data with model predictions favour small values of heavy-quark $D_{s}$, in the range $4.51.5<2\pi D_{s}T_{\mathrm{c}}<$ 4.5 at the pseudocritical temperature $T_{\mathrm{c}}=155$ $\mathrm{MeV}$, in line with other estimates [259, 255, 54, 242, 260]. This interval corresponds to relaxation times for charm quarks smaller than about 10 fm$/c$, thus of the order of or shorter than the QGP lifetime, actually suggesting that charm quarks might get close to reaching kinetic equilibrium in the medium [55, 256]. A recent estimate that combines lattice QCD to the heavy quark effective theory in the limit $M\gg T$ [256] implies a very small thermalization time, at least in the $p\rightarrow 0$ limit; namely of about 1-1.5 fm/c for charm quarks and 3-5 fm/c for bottom which are significantly smaller than those extrapolated from the phenomenological analysis. This work has been developed in non-quenched QCD which is a very important advancement w.r.t. previous studies, however, it is still in the $M\rightarrow\infty$ limit. In Ref. [85, 261] it has been shown that within a quasi-particle model (QPM-Catania) one expects that $D_{s}$ has a significant mass dependence from the charm quark mass to the limit $M\rightarrow\infty$, decreasing by about a factor 1.7. This could to a large extent compensate the discrepancy between the value of $2\pi TD_{s}$ from the lQCD driven estimate w.r.t. the one extrapolated from the phenomenology for charm quarks. However, an even more recent extension of the study in Ref. [256] has done a first extrapolation to finite mass finding a very small dependence of $2\pi TD_{s}$ on $M_{Q}$. This, if confirmed by further studies within lQCD and heavy quark effective field theory, opens the question on how to reconcile them with the phenomenological estimates. At this stage, we only mention that a preliminary study have indicated that an initial glasma phase may lead to infer smaller values of $D_{s}$ [262], and more recently it has been suggested that also memory effect (non-Markovian dynamics) may affect the phenomenological extrapolation of $D_{s}$ [263], especially for charm quarks. Therefore, with more precise data becoming available with the ALICE upgrade (especially at low $p_{\mathrm{T}}$), it is likely that will be possible to achieve a significant advancement in the understanding of the heavy quark dynamics.
@@
However, even within the uncertainties of the mass dependence of the diffusion or drag coefficient, it remains that because of the larger mass, beauty quarks are expected to diffuse less than charm quarks and to have a significantly larger relaxation time that is at least a factor of two [85, 263] if not a factor of three larger (if $D_{S}$ is nearly mass independent) as suggested by the ratio $M_{b}/M_{c}$. Energy loss is also expected to be smaller for beauty than charm quarks. In line with these expectations, experimental measurements of beauty hadrons or their decay products show higher $R_{\mathrm{AA}}$ and smaller $v_{2}$ with respect to charm or light-flavour hadrons [264, 224].
@@
The hadronization schemes used in transport models have a significant impact on the predicted values of $R_{\mathrm{AA}}$ and $v_{2}$ and on their dependence on $p_{\mathrm{T}}$. The effect of hadronization via recombination on D-meson $R_{\mathrm{AA}}$ and $v_{2}$ is illustrated in Fig. 23 where ALICE measurements in Pb–Pb collisions in the centrality intervals 0–10% ($R_{\mathrm{AA}}$, left panel) and 30–50% ($v_{\mathrm{2}}$, right panel) are compared with the expectations of the PHSD, POWLANG, and Catania models with and without the inclusion of recombination [232]. In a medium that is radially expanding with a common velocity, hadronization via quark recombination tends to equilibrate particle velocities towards the medium expansion velocity. The combined effect of recombination, diffusion, and energy loss on the $R_{\mathrm{AA}}$ is the appearance of a maximum at momenta not much larger than the hadron mass. For what concerns the elliptic flow, the $v_{\mathrm{2}}$ of heavy-flavour hadrons formed via recombination derives from a combination of the charm and light quarks flows. From the comparison of model expectations with and without hadronization via recombination, shown in the right panel of Fig. 23, it can be appreciated the effect of recombination on charm-meson $v_{\mathrm{2}}$ in increasing $v_{\mathrm{2}}$ and pushing its maximum at higher momenta. Recombination is an essential ingredient for reproducing the data [267, 67, 221, 255, 101].
@@
All models that can reproduce the large observed $v_{2}$ of D mesons [237] predict that it partly derives from hadronization via coalescence. In order to single out the contribution of hadronization from the contribution of charm-quark diffusion in the generation of charm flow, the modification in Pb–Pb collisions with respect to pp collisions of the $p_{\mathrm{T}}$-differential yield ratios of different heavy-flavour particle species must be studied. In particular, models based on statistical hadronization or coalescence were initially designed to describe hadron formation in AA collisions. They were expecting an increase of the production of heavy-flavour baryons relative to heavy-flavour mesons and of hadrons with strange quarks relative to non-strange ones in heavy-ion with respect to pp collisions, the latter originally thought to deliver systems with properties closer to $\mathrm{e^{+}e^{-}}$ collisions and in which “in-vacuum” fragmentation should have been the dominant hadronization mechanism. Such an expectation was motivated by the standard lore that only the large reservoir of partons and high energy densities of nucleus–nucleus collision systems could make it likely for quark coalescence to become a relevant process compared to the mechanism of fragmentation, which requires the creation of quarks from the vacuum. Hence, the typical signature of coalescence, that is, an enhancement of baryon-to-meson ratios, was not expected to occur in pp collisions. However, as discussed in the previous sections, the data indicates that, particularly in the heavy-flavor sector, pp collisions exhibit significant differences compared to smaller collision systems such as $\mathrm{e^{+}e^{-}}$ and $\mathrm{e^{\pm}p}$. These differences result in modifications to the particle ratios, providing a sensitive test of changes in the hadronization dynamics. The coalescence plus fragmentation approach, such as the Catania model, suggests that even if pp collisions result in the formation of a QGP droplet, D meson production is likely still dominated by the fragmentation process. Meanwhile, the coalescence process becomes dominant for baryon production, significantly increasing it in pp collisions compared to $\mathrm{e^{+}e^{-}}$ collisions. This outcome, though initially unexpected, naturally follows from the application of coalescence dynamics and effectively explains the large increase in the measured $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio. However, the enhancement in pp collisions can also be explained in the context of PYTHIA, but, also in this case, the main modification affects the baryon production that has to come from colour reconnections of quarks and gluons in topologies irrelevant in $\mathrm{e^{+}e^{-}}$ collisions but emerging with an increasing number of MPI, hence particularly in high-multiplicity events, in hadronic collisions. This also shows that already in pp (in particular for high-multiplicity events) more local recombination of quarks (minimizing string energy) becomes quite relevant for baryon production. The same coalescence modelling shows that in AA collisions the large-size reservoir of high-density quarks makes a further difference making the recombination the dominant mechanism for both heavy quark mesons and baryons.
@@
The $\Lambda_{\mathrm{c}}^{+}$ is the only heavy-flavour baryon whose production has been measured in nucleus–nucleus collisions. The left panel of Fig. 25 shows the evolution of the $p_{\mathrm{T}}$-integrated $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio as a function of the charged-particle multiplicity at midrapidity measured by ALICE in pp, p–Pb, and Pb–Pb collisions and by STAR in Au–Au collisions [72]. The ratio is approximately 0.5, thus much higher than the $\approx 0.11$ value measured in $\mathrm{e^{+}e^{-}}$ and ep collisions, and, within uncertainties, it does not show a dependence on multiplicity. The same ratio in pp (minimum bias) and in Pb–Pb collisions is expected by the Catania and TAMU models, which reproduce also quantitatively the data. Also the SHMc model does not expect a dependence on centrality in Pb–Pb collisions. However, it predicts a ratio around 0.2 and underestimates the data. It must be noted that this version of the SHMc does not include the augmented set of high-mass baryon states considered by TAMU. In the SHMc charm is not in chemical equilibrium and the abundance of charm hadron species depends on the charm fugacity factor, which however cancels in single-charm particle ratios, and by the chemical freezeout temperature, which results in being independent of centrality. Similarly, in the TAMU, though the hadronization process is modelled differently in pp and Pb–Pb collisions, particle ratios mainly depend on statistical factors. For the Catania model, the similarity between pp and Pb–Pb values derives from the similar description of the bulk of light-flavour quarks as a QGP medium with similar temperatures in the two systems and from the applied normalization condition that coalescence probability equals unity at $p_{\mathrm{T}}^{\mathrm{hadron}}=0$. There is still a dependence on the volume of the hadronizing fireball and a subdominat dependence on the radial flow but the last has a weak effect on the $p_{T}$-integrated yield. In pp collisions, the PYTHIA 8 Monash tune is even lower than the $\mathrm{e^{+}e^{-}}$ expectation and largely independent of event multiplicity. Contrastingly, PYTHIA 8 with CR-BLC expects a significance increase of the ratio with multiplicity, deriving from the larger probability of forming junction colour connections in events with a higher number of MPI, which is not supported by the data. The region of very low multiplicities is not covered by measurements, thus leaving open the possibility that $\mathrm{e^{+}e^{-}}$ values are recovered when the hadronic activity in the event is very small, in line with what the $\Lambda_{\mathrm{b}}^{0}/\mathrm{B}^{0}$ ratio at forward rapidity, discussed in Sect. III.4, suggests. It is worth mentioning, that the average multiplicity in $\mathrm{e^{+}e^{-}}$ collisions at LEP energies, where most particles are part of high-energy dijets, does not differ much from the average multiplicity in minimum bias pp collisions at LHC energies (see Fig. 19.6 of Ref. [49]). This suggests that the number of produced particles is not the only relevant parameter and rather the whole event topology and the impact of the underlying event and MPI in hadronic collisions, even at the lowest multiplicity, must be taken into account for explaining the differences observed with respect to $\mathrm{e^{+}e^{-}}$ collisions.

### v3

@@ III.4 Measurements as a function of multiplicity
The LHCb Collaboration recently released the measurement of the $\Lambda_{\mathrm{b}}^{0}/{\mathrm{B}^{0}}$ cross-section ratio as a function of $p_{\mathrm{T}}$ and multiplicity [14]. The $p_{\mathrm{T}}$-integrated ratio is shown as a function of relative multiplicity in Fig. 20. In the lowest multiplicity interval, the LHCb data approaches the baryon fraction measured in $\mathrm{e^{+}e^{-}}$ collisions at the LEP. The distinct rise of the baryon fraction with multiplicity reaches a plateau at $\sim$ 0.5 for collisions that produce more than twice the average number of VELO tracks.
These last two measurements show that b-quark hadronization is different in $\mathrm{e^{+}e^{-}}$ and pp collisions, thus it is not an universal process. This may provide indication that quark coalescence plays an important role in forming beauty hadrons at the LHC, or that one should more carefully examine string models with multiplicity dependencies. The trend observed for the $p_{\mathrm{T}}$-integrated $\Lambda_{\mathrm{b}}^{0}/{\mathrm{B}^{0}}$ ratio resembles a “turn-on” curve which may mark the transition from two different regimes, with the onset of a different hadronization mechanism in pp collisions becoming very quickly effective as soon as a minimal hadronic activity is present. It would be interesting to try correlating the observed trend with the number of MPI. A work, in which a set of unobserved excited beauty quark states are included, seemed to be able to qualitatively explain the system size dependence of these ratios in terms of canonical suppression of baryons and strangeness toward smaller multiplicities [217]. It must be noted that the evolution observed for the $p_{\mathrm{T}}$-integrated $\Lambda_{\mathrm{b}}^{0}/{\mathrm{B}^{0}}$ ratio does not contrast what was previously discussed about the evolution with multiplicity of the $p_{\mathrm{T}}$-differential $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio measured by ALICE at midrapidity. This will be further discussed in Sect. III.6.
@@ III.5 Proton–nucleus collisions
The LHCb collaboration measured the $\Lambda_{\mathrm{c}}^{+}$ and $\Xi_{\mathrm{c}}^{+}$ production and their ratios to $\mathrm{D^{0}}$ at forward and backward rapidity in p–Pb collisions [26, 27]. When compared with the measurements at midrapidity [25, 17], the baryon-to-meson ratios are lower for $p_{\mathrm{T}}$ $<$ 8 GeV/$c$, while at higher $p_{\mathrm{T}}$ the measurements are consistent within uncertainties. The origin of this tension is not yet fully understood, for a more detailed discussion on this open point see Sect. IV. Also the ratio of $\mathrm{D_{s}^{+}}$ production to that of $\mathrm{D^{0}}$ and $\mathrm{D^{+}}$ was studied in p–Pb collisions by ALICE [231] and LHCb [233]. Without selections on multiplicity, the ratios are observed to be consistent with those measured in pp collisions. On the other hand, while ALICE does not observe a significant dependence on centrality, LHCb reports an increase of the ratio with multiplicity, with consistent values at forward and backward rapidity. However, the data uncertainty prevent conclusions on a possible dependence on rapidity of the ratios, as well as whether LHCb values converge to the values measured by ALICE in Pb–Pb collisions [234, 235].
@@ III.6 Nucleus–nucleus collisions
Several measurements in Au–Au collisions at RHIC and Pb–Pb collisions at the LHC demonstrate that charm and beauty quarks interact strongly with the QGP constituents [239, 240, 241, 242, 243, 178, 244, 245, 246, 247, 248, 249, 235, 250, 251, 252, 253, 254, 255, 256, 257, 258, 259, 260, 261, 176, 262, 263, 264, 265, 266]. The $R_{\mathrm{AA}}$ values observed for both charm and beauty hadrons are significantly smaller than unity at intermediate $p_{\mathrm{T}}$ and larger for beauty than charm, indicating a significant energy loss of both quarks in the medium and evidencing the expected dependence of energy loss on the quark mass. At moderate/low $p_{\mathrm{T}}$ the $v_{2}$ of prompt D mesons is large, close to that observed for charged pions [251]. This can indicate that charm quarks are strongly coupled to the system and are pushed towards kinetic equilibrium via rescattering with the medium constituents. It can also signal coalescence as a hadronization process relevant to charm quarks through which charm hadrons partly inherit the flow of light quarks, a relevant aspect that could be more deeply investigated through heavy-light correlation of the anisotropic flow $v_{n}$ [267, 99]. At low $p_{\mathrm{T}}$, where elastic processes are the most important ones, heavy quark transport in the medium can be described as a diffusion process [268, 269, 74, 270]. The main transport parameter is the spatial diffusion coefficient $\mathcal{D}_{s}$, which is related to the kinetic equilibration time of a heavy-quark with mass $m_{\mathrm{Q}}$, $\tau_{\mathrm{Q}}=\mathcal{D}_{s}m_{\mathrm{Q}}/T$ [271, 272]. The comparison of $R_{\mathrm{AA}}$ and $v_{\mathrm{2}}$ data with model predictions favour small values of heavy-quark $\mathcal{D}_{s}$, in the range $4.51.5<2\pi\mathcal{D}_{s}T_{\mathrm{c}}<$ 4.5 at the pseudocritical temperature $T_{\mathrm{c}}=155$ $\mathrm{MeV}$, in line with other estimates [273, 269, 73, 256, 274]. This interval corresponds to relaxation times for charm quarks smaller than about 10 fm$/c$, thus of the order of or shorter than the QGP lifetime, actually suggesting that charm quarks might get close to reaching kinetic equilibrium in the medium [74, 270]. A recent estimate that combines lattice QCD to the heavy quark effective theory in the limit $M\gg T$ [270] implies a very small thermalization time, at least in the $p\rightarrow 0$ limit; namely of about 1-1.5 fm/c for charm quarks and 3-5 fm/c for bottom which are significantly smaller than those extrapolated from the phenomenological analysis. This work has been developed in non-quenched QCD which is a very important advancement w.r.t. previous studies, however, it is still in the $M\rightarrow\infty$ limit. In Ref. [100, 119] it has been shown that within a quasi-particle model (QPM-Catania) one expects that $\mathcal{D}_{s}$ has a significant mass dependence from the charm quark mass to the limit $M\rightarrow\infty$, decreasing by about a factor 1.7. This could to a large extent compensate the discrepancy between the value of $2\pi T\mathcal{D}_{s}$ from the lQCD driven estimate w.r.t. the one extrapolated from the phenomenology for charm quarks. However, an even more recent extension of the study in Ref. [270] has done a first extrapolation to finite mass finding a very small dependence of $2\pi T\mathcal{D}_{s}$ on $M_{Q}$. This, if confirmed by further studies within lQCD and heavy quark effective field theory, opens the question on how to reconcile them with the phenomenological estimates. At this stage, we only mention that a preliminary study have indicated that an initial glasma phase may lead to infer smaller values of $\mathcal{D}_{s}$ [275], and more recently it has been suggested that also memory effect (non-Markovian dynamics) may affect the phenomenological extrapolation of $\mathcal{D}_{s}$ [276], especially for charm quarks. Therefore, with more precise data becoming available with the ALICE upgrade (especially at low $p_{\mathrm{T}}$), it will be possible to achieve a significant advancement in the understanding of the heavy quark dynamics.
@@
However, even within the uncertainties of the mass dependence of the diffusion or drag coefficient, it remains that because of the larger mass, beauty quarks are expected to diffuse less than charm quarks and to have a significantly larger relaxation time that is at least a factor of two [100, 276] if not a factor of three larger (if $\mathcal{D}_{s}$ is nearly mass independent) as suggested by the ratio $M_{b}/M_{c}$. Energy loss is also expected to be smaller for beauty than charm quarks. In line with these expectations, experimental measurements of beauty hadrons or their decay products show higher $R_{\mathrm{AA}}$ and smaller $v_{2}$ with respect to charm or light-flavour hadrons [277, 239].
@@
The hadronization schemes used in transport models have a significant impact on the predicted values of $R_{\mathrm{AA}}$ and $v_{2}$ and on their dependence on $p_{\mathrm{T}}$. The effect of hadronization via recombination on D-meson $R_{\mathrm{AA}}$ and $v_{2}$ is illustrated in Fig. 23 where ALICE measurements in Pb–Pb collisions in the centrality intervals 0–10% ($R_{\mathrm{AA}}$, left panel) and 30–50% ($v_{\mathrm{2}}$, right panel) are compared with the expectations of the PHSD, POWLANG, and Catania models with and without the inclusion of recombination [247]. In a medium that is radially expanding with a common velocity, hadronization via quark recombination tends to equilibrate particle velocities towards the medium expansion velocity. The combined effect of recombination, diffusion, and energy loss on the $R_{\mathrm{AA}}$ is the appearance of a maximum at momenta not much larger than the hadron mass. For what concerns the elliptic flow, the $v_{\mathrm{2}}$ of heavy-flavour hadrons formed via recombination derives from a combination of the charm and light quarks flows. From the comparison of model expectations with and without hadronization via recombination, shown in the right panel of Fig. 23, it can be appreciated the effect of recombination on charm-meson $v_{\mathrm{2}}$ in increasing $v_{\mathrm{2}}$ and pushing its maximum to higher momenta. Recombination is an essential ingredient for reproducing the data [280, 86, 236, 269, 116].
@@
All models that can reproduce the large observed $v_{2}$ of D mesons [251] predict that it partly derives from hadronization via coalescence. In order to single out the contribution of hadronization from the contribution of charm-quark diffusion in the generation of charm flow, the modification in Pb–Pb collisions with respect to pp collisions of the $p_{\mathrm{T}}$-differential yield ratios of different heavy-flavour particle species must be studied. In particular, models based on statistical hadronization or coalescence were initially designed to describe hadron formation in AA collisions. They were expecting an increase of the production of heavy-flavour baryons relative to heavy-flavour mesons and of hadrons with strange quarks relative to non-strange ones in heavy-ion with respect to pp collisions, the latter originally thought to deliver systems with properties closer to $\mathrm{e^{+}e^{-}}$ collisions and in which “in-vacuum” fragmentation should have been the dominant hadronization mechanism. Such an expectation was motivated by the standard lore that only the large reservoir of partons and high energy densities of nucleus–nucleus collision systems could make it likely for quark coalescence to become a relevant process compared to the mechanism of fragmentation, which requires the creation of quarks from the vacuum. Hence, the typical signature of coalescence, that is, an enhancement of baryon-to-meson ratios, was not expected to occur in pp collisions. However, as discussed in the previous sections, the data indicates that, particularly in the heavy-flavour sector, pp collisions exhibit significant differences compared to smaller collision systems such as $\mathrm{e^{+}e^{-}}$ and $\mathrm{e^{\pm}p}$. These differences result in modifications to the particle ratios, providing a sensitive test of changes in the hadronization dynamics. The coalescence plus fragmentation approach, such as the Catania model, suggests that even if pp collisions result in the formation of a QGP droplet, D meson production is likely still dominated by the fragmentation process. Meanwhile, the coalescence process becomes dominant for baryon production, significantly increasing it in pp collisions compared to $\mathrm{e^{+}e^{-}}$ collisions. This outcome, though initially unexpected, naturally follows from the application of coalescence dynamics and effectively explains the large increase in the measured $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio. However, the enhancement in pp collisions can also be explained in the context of PYTHIA, but, also in this case, the main modification affects the baryon production that has to come from colour reconnections of quarks and gluons in topologies irrelevant in $\mathrm{e^{+}e^{-}}$ collisions but emerging with an increasing number of MPI, hence particularly in high-multiplicity events, in hadronic collisions. This also shows that already in pp (in particular for high-multiplicity events) more local recombination of quarks (minimizing string energy) becomes quite relevant for baryon production. The same coalescence modelling shows that in AA collisions the large-size reservoir of high-density quarks makes a further difference making the recombination the dominant mechanism for both heavy quark mesons and baryons.
@@
The $\Lambda_{\mathrm{c}}^{+}$ is the only heavy-flavour baryon whose production has been measured in nucleus–nucleus collisions. The left panel of Fig. 25 shows the evolution of the $p_{\mathrm{T}}$-integrated $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio as a function of the charged-particle multiplicity at midrapidity measured by ALICE in pp, p–Pb, and Pb–Pb collisions and by STAR in Au–Au collisions [28]. The ratio is approximately 0.5, thus much higher than the $\approx 0.11$ value measured in $\mathrm{e^{+}e^{-}}$ and ep collisions, and, within uncertainties, it does not show a dependence on multiplicity. The same ratio in pp (minimum bias) and in Pb–Pb collisions is expected by the Catania and TAMU models, which reproduce also quantitatively the data. Also the SHMc model does not expect a dependence on centrality in Pb–Pb collisions. However, it predicts a ratio around 0.2 and underestimates the data. It must be noted that this version of the SHMc does not include the augmented set of high-mass baryon states considered by TAMU. In the SHMc charm is not in chemical equilibrium and the abundance of charm hadron species depends on the charm fugacity factor, which however cancels in single-charm particle ratios, and by the chemical freezeout temperature, which results in being independent of centrality. Similarly, in the TAMU, though the hadronization process is modelled differently in pp and Pb–Pb collisions, particle ratios mainly depend on statistical factors. For the Catania model, the similarity between pp and Pb–Pb values derives from the similar description of the bulk of light-flavour quarks as a QGP medium with similar temperatures in the two systems and from the applied normalization condition that coalescence probability equals unity at $p_{\mathrm{T}}^{\mathrm{hadron}}=0$. There is still a dependence on the volume of the hadronizing fireball and a subdominat dependence on the radial flow but the last has a weak effect on the $p_{T}$-integrated yield. In pp collisions, the PYTHIA 8 Monash tune is even lower than the $\mathrm{e^{+}e^{-}}$ expectation and largely independent of event multiplicity. Contrastingly, PYTHIA 8 with CR-BLC expects a significance increase of the ratio with multiplicity, deriving from the larger probability of forming junction colour connections in events with a higher number of MPI, which is not supported by the data. The region of very low multiplicities is not covered by measurements, thus leaving open the possibility that $\mathrm{e^{+}e^{-}}$ values are recovered when the hadronic activity in the event is very small, in line with what the $\Lambda_{\mathrm{b}}^{0}/\mathrm{B}^{0}$ ratio at forward rapidity, discussed in Sect. III.4, suggests. It is worth recalling, that, as discussed in Sect. III.4, the average multiplicity in $\mathrm{e^{+}e^{-}}$ collisions at LEP energies, where most particles are part of high-energy dijets, may not differ tremendously from the average multiplicity in minimum bias pp collisions at LHC energies. This suggests that the number of produced particles is not the only relevant parameter and rather the whole event topology and the impact of the underlying event and MPI in hadronic collisions, even at the lowest multiplicity, must be taken into account for explaining the differences observed with respect to $\mathrm{e^{+}e^{-}}$ collisions.